# The polynomial set associated with a fixed number of matrix-matrix multiplications

### v1

@@ 4.1 Minimality of unreduced evaluation schemes
As a consequence of the equivalence theorems in the previous section, we can conclude that any unreduced evaluation scheme can be be assumed to have the form
(61) $[A|B]=\left[\begin{array}[]{ccccccc|cccccccc}0&1&&&&&&0&1&&&&\\ 0&a&1&&&&&0&0&1&&&\\ 0&a&a_{33}&1&&&&0&b&{a_{33}+1}&1&&\\ 0&a&a&a&1&&&0&b&b&b&1&\\ \vdots&\vdots&\vdots&\vdots&\ddots&\ddots&&\vdots&\vdots&\vdots&\vdots&\ddots&\ddots\\ 0&a&a&a&\cdots&a&1&0&b&b&b&\cdots&b&1\\ \end{array}\right],$
with an arbitrary $c$-vector. The dimension of $\Pi_{2^{m}}^{*}$ is determined from the rank of the Jacobian for a generic element in $\Pi_{2^{m}}^{*}$. A generic element in $\Pi_{2^{m}}^{*}$ is unreduced. Hence, we can conclude that the dimension is bounded by the number of free variables. The parameterization (61) contains $m^{2}$ parameters, and therefore, similar to (12), we have the following.
Corollary 4.1.
@@
In order to determine the sharpness of the bound, we have done numerical computations in high precision arithmetic. In our simulations, the numerical rank of the the Jacobian (as defined in Section 2.2) for random $(A,B,c)$ has always coincided with the bound; verified up until $m=10$. Therefore, this is seemingly a minimal parameterization of unreduced schemes in $\Pi_{2^{m}}^{*}$; to our knowledge the first minimal parameterization.
Conjecture 4.2.
@@
$\dim(\Pi^{*}_{2^{m}})=m^{2}.$
@@ 4.2 Reduced evaluation schemes
In practice, the unreduced evaluation schemes are rarely useful to evaluate a given polynomial of a high degree $d$, due to a lack of degrees of freedom in $\Pi_{2^{m}}^{*}$ for large $m$, in comparison to $\dim(\Pi_{2^{m}})=d+1=2^{m}+1$. If the Hessenberg matrices $A$ and $B$ are reduced, we obtain output polynomials of lower degree, for example the pair
(62) $[A|B]=\left[\begin{array}[]{ccccccc|cccccccc}0&1&&&&&&0&1&&&&\\ 0&\times&1&&&&&0&0&1&&&\\ 0&\times&\times&1&&&&0&\times&\times&1&&\\ 0&\times&\times&\times&1&&&0&\times&\times&1&0&\\ 0&\times&\times&\times&\times&1&&0&\times&\times&\times&1&0\\ 0&\times&\times&\times&\times&\times&1&0&\times&\times&\times&\times&1&0\\ \end{array}\right]$
corresponds to $m=6$ multiplications, but results in a polynomial of degree $32$. With a single reduction, we mean that, in one row we set the last non-zero elements of to zero in either $A$ or $B$. With $r$ reductions we mean the repeated application of a single reduction. Note that we can still normalize each row since the transformation theorems also applies to reduced systems. Hence, we lose a degree of freedom with every reduction, and due to Corollary 4.1, we expect the corresponding dimension to be
@@
In the follow section we proceed by studying reduced evaluation schemes, that lead to specific polynomial degrees, and describe ways to compute $(A,B,c)$ for that reduction structure, for a polynomial given in a monomial basis.
@@
Figure 1: The polynomial degrees when the Hessenberg matrices in the triplet $(A,B,c)$ are reduced Hessenberg matrices. The parameter $\epsilon=0.1$ is selected for visualization reasons. Any point above the blue curve can be discarded to not completely contain the corresponding polynomial subset, in the sense of (64), as it has an insufficient number of degrees of freedom.
@@ 5 Polynomial subsets
This section is devoted to the study of the question: What is largest $d$ such that all $d$-degree polynomials can be computed with $m$ multiplications? Formally, we use two versions of this problem
@@
We consider the left-hand side when possible, and otherwise study the right-hand side in order to avoid limit cases that can occur similar to (8).
The previous section stressed that we must consider reduced matrices. For a given reduction, we can compute the corresponding polynomial degree. Hence, we can investigate candidate solutions to (64) by considering all possible reductions. This approach is depicted in Figure 1, which illustrates all combinations of reductions for up to $m\leq 7$ multiplications and $r\leq 7$ reductions For instance, (62) corresponds to the scenario (horizontal axis) with $m=6$ multiplications and $r=3$ reductions, and achieves a polynomial degree of $32$, as shown on the vertical axis of the figure. For identifying optimal solutions to (64), higher polynomial degrees are preferable. On the other hand, a too high degree will lead to an insufficient number of degrees of freedom. More precisely, (63) gives a bound on the degree of the admissable candidate solutions to (64) for a given number of multiplications and reductions. This is visualized with a blue line in the figure. Hence, for the purpose of studying (64), we can disregard points above this line.
The problem (64) gets more and more complicated for the higher $m$ is. For $m=3$, the problem is essentially already solved in [24] since an explicit procedure is provided to compute almost all polynomials of degree $8$ with $m=3$ multiplications, i.e.,
@@
For $m=4$, similar constructions are also provided in [24], providing a method for degree $12$. Another approach with smaller number of assumptions on the monomial coefficients is provided in Section 5.1. From Figure 1, we see that this is the highest admissible degree and therefore conclude that this is optimal. For $m\geq 5$ we were not able to solve the problem analytically, and instead resort to computational tools. More precisely, we frame the problem with a given reduction and structure as a system of polynomial equations. For $m=5$ we use the package HomotopyContinuation.jl [4], to find seemingly all solutions. For $m=6$ and $m=7$, the system was too complicated, and we were not able to find solutions with this package. However, by using the software [14], we were able to construct locally convergent iterative methods and in all specified test cases, able to find solutions. These simulations combined with reasoning with admissible degrees in Figure 1, we conclude conjectures concerning the solution to (64). For reproducibility, all simulations (including starting values) are provided in the publicly available github repository: https://github.com/GustafLorentzon/polynomial-set-paper.
@@ 5.1 Four multiplications
When we study $m=4$, we identify from Figure 1 that the highest degree of the admissible polynomials is $12$, since without reductions we only have $16$ degrees of freedom, which cannot parameterize $\Pi_{16}$. The solution to (64) is indeed 12. This can already be concluded from the method in [24, pg 237] which is a method to evaluate polynomials of degree 12, which are given in their monomial basis, using only $m=4$ multiplications. In our terminology this corresponds to the reduction $a_{2,3}=0$, and additionally $a_{4,2}=0$. The method [24, pg 237] involves square roots of expression containing the monomial coefficients, as well as divisions of certain quantities leading some exceptions corresponding to certain limit cases. Therefore from [24] we conclude that we have
@@
We now present a slightly more general method for evaluating any polynomial in $\Pi_{12}$ in four multiplications, that does not involve square roots or divisions except for the leading monomial coefficient. Consider the evaluation schemes with the following structure
@@
Suppose $\alpha_{0},\ldots,\alpha_{12}$ represent a given polynomial $p(X)=\alpha_{0}I+\alpha_{1}X+\cdots+\alpha_{12}X^{12}\in\Pi_{12}$. If we expand the parameterization we obtain a multivariate polynomial system of equations, one for each monomial coefficient in the output polynomial. In the terminology of Section 2.2 the system corresponds to considering the 00th,$\ldots$,$12$th derivatives of the equation $\Phi(A,B,c)(x)=p(x)$ with respect to $x$ evaluated at $x=0$. In this case we have $13$ equations in the $13$ variables. To solve solve this system, we first introduce the auxiliary variables $\beta_{43}=b_{43}+a_{43}$ and $\beta_{42}=b_{42}+a_{42}$. This system is explicitly solvable by considering the equations in the output polynomial ordered in descending degree, such that the equation corresponding to $\alpha_{12}$ is treated first. The system is in this sense triangular, which was also crucial for the construction in [24, Section 3]. The solution to the system is given by following sequence of equations:
@@
With the conditions $c_{2}=\alpha_{1}$, $c_{1}=\alpha_{0}$, $b_{43}=\beta_{43}-a_{43}$ and $b_{42}=\beta_{42}-a_{42}$, we have explictly computed all variables in (66).
Recall that $\alpha_{12}\neq 0$ for $p\in\Pi_{12}$, therefore we have made no assumptions other than the degree of the polynomial. Moreover, the formulas preserve the algebraic structure of the variables, e.g., if $\alpha_{0},\ldots,\alpha_{12}\in\mathbb{R}$, then $(A,B,c)$ is a real triplet, i.e., the evaluation coefficients are real. We conclude that
@@
holds in both a real as well as in a complex sense.
@@ 5.2 Five multiplications
For $m=5$ multiplications we see in Figure 1 that the highest degree of the admissible polynomials is $20$. To our knowledge, state-of-the-art is $d=18$ as given in [25, Equation (17)-(19)] with $s=2$ which is based on [24] combined with the Paterson-Stockmeyer evaluation. In our terminology, that approach corresponds to the reduction $a_{23}=a_{56}=0$ and additionally imposing $a_{42}=0$. The reduction in Figure 1 leading to a polynomial of degree $d=20$ corresponds to $a_{45}=a_{56}=0$.
We were not able to explicitly derive a solution to the multivariate polynomial system for the structure $d=20$ analytically, but instead needed to resort to computational tools. The Julia package HomotopyContinuation.jl [4] includes methods to solve polynomial systems of equations based on numerical continuation and with advanced initialization of starting points for the homotopy method. We implemented the system equations stemming from the considering each monomial coefficient for the data structures of this package. Since we have more variables than equations for this structure, we empirically fixed some variables, and chose which variable to fix by trying to reduce the total degree of the system as much as possible. We additionally solved those equations explicitly that could be solved analytically, e.g., first and last equations.
With this setup we were able to compute $(A,B,c)$ for a large number of given polynomials $p\in\Pi_{20}$, including the truncated Taylor expansion of the exponential, as well as the function $1/(1-x)$. The simulation as well as code that computes the corresponding matrix functions are given in the github repository, both for Julia and matlab. For conciseness we report only the numbers for the matrix exponential in the following.
Due numerical condition, we prefer solutions that avoid very large and very small numbers. In our case, to mitigate large numbers we used a scaling of the input. For the exponential we approximate $e^{\alpha x}$ with $\alpha=8$ fixed, since this makes $c_{7}=\alpha^{20}/20!\approx 0.47$ in the order of magnitude one. Note that the scaling of the input can be compensated in the table, although this was not deemed as numerically useful, and hence not used below.
HomotopyConinuation.jl found several solutions and the solution with the smallest values was the following. The parameters were

### v3

@@ 4.1 Minimality of unreduced evaluation schemes
As a consequence of the equivalence theorems in the previous section, we can conclude that any unreduced evaluation scheme can be assumed to have the form
(56) $A|B]=\left[\begin{array}[]{ccccccc|cccccccc}0&1&&&&&&0&1&&&&\\ 0&a&1&&&&&0&0&1&&&\\ 0&a&a_{33}&1&&&&0&b&{a_{33}+1}&1&&\\ 0&a&a&a&1&&&0&b&b&b&1&\\ \vdots&\vdots&\vdots&\vdots&\ddots&\ddots&&\vdots&\vdots&\vdots&\vdots&\ddots&\ddots\\ 0&a&a&a&\cdots&a&1&0&b&b&b&\cdots&b&1\\ \end{array}\right],$
with an arbitrary $c$-vector. The dimension of $\Pi_{2^{m}}^{*}$ is determined from the rank of the Jacobian for a generic element in $\Pi_{2^{m}}^{*}$. A generic element in $\Pi_{2^{m}}^{*}$ is unreduced; hence, the dimension is bounded by the number of free variables in a parameterization of the unreduced schemes. The parameterization (56) contains $m^{2}$ parameters and, therefore, similar to (7), we have the following.
Corollary 9.
@@
It turns out that this bound is sharp; i.e., we can establish a lower bound which is also $m^{2}$. To this end it is sufficient to find a triplet $(A,B,c)$ such that the Jacobian of the output $p(X)$ with respect to the free variables has rank $m^{2}$. Many trivial choices of unreduced $(A,B,c)$, e.g., a triplet corresponding to only a sequence of squaring operations, lead to a points in $\Pi_{2^{m}}^{*}$ whose rank is less than $m^{2}$. The triplet with a full Jacobian rank and leading to the simplest formulas for the Jacobian that we could find is the following:
(57) $A|B]=\left[\begin{array}[]{cccccc|cccccc}0&1&&&&&0&1&&&&\\ 0&1&1&&&&0&1&1&&&\\ 0&1&0&1&&&0&1&1&1&&\\ \vdots&\vdots&\vdots&\ddots&\ddots&&\vdots&\vdots&\vdots&\ddots&\ddots&\\ 0&1&0&\cdots&0&1&0&1&1&\cdots&1&1\\ \end{array}\right]$
and $c=[0,\ldots,0,1]$.
In order to bound the rank, it is sufficient to find $m^{2}$ linearly independent columns in the Jacobian. A sufficient condition for linear independence for polynomials is that they have distinct degrees. Based on that reasoning, we now establish linear combinations of partial derivatives of the output of the evaluation, i.e., $p(X)$, with respect to the elements of $A,B$ and $c$. We obtain $m^{2}$ distinct degrees.
Although the choice (57) leads to the simplest derivation we could establish analytically, it is admittedly not very simple. The proof of the general case can be found in the appendix. For illustration, we sketch the derivation for $m=4$, specifying the partial derivatives, which can be computed with symbolic computation tool (e.g., those described in the next section). In the full proof in the appendix we provide the details without such tools. We base the proof, as well as this sketch on a separation into cases, each case leading to a polynomial degrees that are distinct, and adding up to $m^{2}=16$ different degrees.
Case 1 corresponds to forming derivatives with respect to elements of $c$: $\frac{\partial p}{\partial c_{1}}=1,\ \frac{\partial p}{\partial c_{2}}=X,\ \frac{\partial p}{\partial c_{3}}=X^{2},\ \frac{\partial p}{\partial c_{4}}=Q_{4},\ \frac{\partial p}{\partial c_{5}}=Q_{5},\ \frac{\partial p}{\partial c_{6}}=Q_{6}$. Since $(A,B,c)$ in (57) is unreduced, $Q_{4},Q_{5}$ and $Q_{6}$ have maximal degree and we have that
(58) $\deg\left(\frac{\partial p}{\partial c_{i}}\right)=\deg(Q_{i})=2^{i-2},i=2,\ldots,m+2.$
Case 2 corresponds to forming derivatives with respect to $a_{i,j}$. For example, we have
(59) $\frac{\partial p}{\partial a_{4,4}}$ $=$ $X^{12}+6X^{11}+16X^{10}+26X^{9}+\mathcal{O}(X^{8})$
(60) $\frac{\partial p}{\partial a_{4,2}}$ $=$ $X^{9}+4X^{8}+7X^{7}+8X^{6}+7X^{5}+\mathcal{O}(X^{4}).$
As shown in the proof of the theorem, the general formula for the degrees in Case 2 is
(61) $\deg\left(\frac{\partial p}{\partial a_{i,j}}\right)=2^{m}-2^{i-1}+2^{j-2},\textrm{ for }i=2,\ldots,m,\;\;j=2,\ldots,i$
Case 3 stems from the observation that we obtain the same degree if we differentiate with respect to $a_{i,j}$ and $b_{i,j}$. Therefore we form the difference in order to obtain a different degree. For example,
(62) $\frac{\partial p}{\partial b_{4,2}}=X^{9}+4X^{8}+7X^{7}+8X^{6}+6X^{5}+\mathcal{O}(X^{4})$
and the difference with (60) is
(63) $\frac{\partial p}{\partial b_{4,2}}-\frac{\partial p}{\partial a_{4,2}}=-X^{5}-2X^{4}-2X^{3}.$
The degree $5$ is distinct from the degrees in Cases 1 and 2. In the appendix we prove that the general formula for the degrees in Case 3 is
(64) $\deg\left(\frac{\partial p}{\partial b_{i,j}}-\frac{\partial p}{\partial a_{i,j}}\right)=2^{m}-2^{i}+2^{i-2}+2^{j-2},\textrm{ for }i=3,\ldots,m,\;\;j=2,\ldots,i-1$
and that they are distinct from previous cases.
Case 4 is based on the observation that if we use the idea from Case 3 for $j=i=m$, we get a degree which already included in Case 1. In order to establish a degree not present in the previous cases, we must form a linear combination of partial derivatives with respect to the elements $c_{m+1}$, $a_{m,m}$, $b_{m,m}$, $a_{m,2}$ and $b_{m,2}$. Consider the following identities:
(65) $\frac{\partial p}{\partial b_{4,4}}$ $=$ $X^{12}+6X^{11}+16X^{10}+26X^{9}+\mathcal{O}(X^{8})$
(66) $\frac{\partial p}{\partial b_{4,4}}-\frac{\partial p}{\partial a_{4,4}}$ $=$ $-X^{8}-4X^{7}-7X^{6}-6X^{5}+\mathcal{O}(X^{4})$
(67) $\frac{\partial p}{\partial c_{5}}$ $=$ $X^{8}+4X^{7}+7X^{6}+8X^{5}+\mathcal{O}(X^{4}).$
By forming the sum of equations (66) and (67), we obtain
$\frac{\partial p}{\partial b_{4,4}}-\frac{\partial p}{\partial a_{4,4}}+\frac{\partial p}{\partial c_{5}}=2X^{5}+4X^{4}+3X^{3}+X^{2}.$
Noting from (63) that this degree coincides with Case 3 for $i=4$ and $j=2$, and we can reduce the degree by forming the sum
$\frac{\partial p}{\partial b_{4,4}}-\frac{\partial p}{\partial a_{4,4}}+\frac{\partial p}{\partial c_{5}}+2\left(\frac{\partial p}{\partial b_{4,2}}-\frac{\partial p}{\partial a_{4,2}}\right)=-X^{3}+X^{2}.$
The resulting degree 3 is distinct from those in Cases 1-3.
From the reasoning above, in this example, we have 6 distinct degrees from Case 1, 6 distinct degrees from Case 2, 3 distinct degrees from Case 3 and 1 degree from Case 4, which yield a total of $m^{2}=16$ distinct degrees.
In the general case, i.e., when $m>4$, we also provide a Case 5, leading to an additional to m4m-4 distinct degrees. The idea for this case is very similar to that of Case 4, and corresponds to forming a linear combination of partial derivatives with respect to the elements $a_{i,i}$, $b_{i,i}$ $a_{i+1}$, $a_{i,2}$ and $b_{i,2}$ for $i<m$.
Theorem 10.
@@
(68) $\dim(\Pi^{*}_{2^{m}})=m^{2}.$
Proof.
See Section A. ∎
As a consequence of Theorem 10, the parameterization (56) is minimal. To our knowledge, this is the first minimal parameterization of unreduced schemes in $\Pi_{2^{m}}^{*}$.
@@ 4.2 Reduced evaluation schemes
In practice, unreduced evaluation schemes are rarely useful for evaluating a given polynomial of high degree $d$, because of the limited number of degrees of freedom in $\Pi_{2^{m}}^{*}$ for large $m$, in comparison to $\dim(\Pi_{2^{m}})=d+1=2^{m}+1$. If the Hessenberg matrices $A$ and $B$ are reduced, we obtain output polynomials of lower degree. For example the pair
(69) $A|B]=\left[\begin{array}[]{ccccccc|cccccccc}0&1&&&&&&0&1&&&&\\ 0&\times&1&&&&&0&0&1&&&\\ 0&\times&\times&1&&&&0&\times&\times&1&&\\ 0&\times&\times&\times&1&&&0&\times&\times&1&0&\\ 0&\times&\times&\times&\times&1&&0&\times&\times&\times&1&0\\ 0&\times&\times&\times&\times&\times&1&0&\times&\times&\times&\times&1&0\\ \end{array}\right]$
corresponds to $m=6$ multiplications but results in a polynomial of degree $32$. By a single reduction, we mean setting the last nonzero element in a row to zero in either $A$ or $B$. With $r$ reductions we mean the repeated application of a single reduction. Note that we can still normalize each row since the transformation theorems are also applicable to reduced systems. Hence, we lose one degree of freedom with every reduction, and due to Corollary 9, we expect the corresponding dimension to be
@@
In the following section we proceed by studying reduced evaluation schemes. More precisely, we study reduced evaluation schemes that lead to specific polynomial degrees and describe ways to compute $(A,B,c)$ for that reduction structure for a polynomial given in a monomial basis.
@@
Figure 1: The polynomial degrees when the Hessenberg matrices in the triplet $(A,B,c)$ are reduced Hessenberg matrices. The parameter $\epsilon=0.1$ is selected for visualization purposes. Any point above the blue curve can be discarded to not completely contain the corresponding polynomial subset, in the sense of (71), due to an insufficient number of degrees of freedom.
@@ 5 Polynomial subsets
This section is devoted to the study of the question: What is the largest $d$ such that all $d$-degree polynomials can be computed with $m$ multiplications? Formally, we use two versions of this problem
@@
We consider the left-hand side when possible, and otherwise study the right-hand side in order to avoid limit cases similar to (5).
The previous section stressed the use of reduced matrices. For a given reduction, we can compute the corresponding polynomial degree. Hence, we can investigate candidate solutions to (71) by considering all possible reductions. This approach is depicted in Figure 1, which illustrates all combinations of reductions for up to $m\leq 7$ multiplications and $r\leq 7$ reductions. For instance, (69) corresponds to the scenario (horizontal axis) with $m=6$ multiplications and $r=3$ reductions and achieves a polynomial degree of $32$, as shown on the vertical axis of the figure.
To identify optimal solutions to (71), higher polynomial degrees are advantageous. However, excessively high degrees may result in insufficient degrees of freedom. More precisely, (70) gives a bound on the degree of the admissible candidate solutions to (71) for a given number of multiplications and reductions. This is visualized with a blue line in the figure. Hence, for the purpose of studying (71) we can disregard points above this line.
The problem (71) becomes increasingly complex as $m$ increases. For $m=3$, the problem is essentially already solved in [24] since an explicit procedure is provided to compute almost all polynomials of degree $8$ with $m=3$ multiplications, i.e.,
@@
For $m=4$, similar constructions are also provided in [24], yielding a method for degree $12$. An alternate approach requiring fewer assumptions on the monomial coefficients is given in Section 5.1. From Figure 1, we see that this is the highest admissible degree and therefore conclude that it is optimal.
For $m\geq 5$ we were not able to solve the problem analytically and instead resorted to computational tools. More precisely, we frame the problem with a given reduction and structure as a system of polynomial equations. For $m=5$, we use the package HomotopyContinuation.jl [4] to find, as far as we can tell, all solutions. For $m=6$ and $m=7$, the system was too complicated, and we were not able to find solutions with this package. However, by using the software [14], we were able to construct locally convergent iterative methods and successfully found solutions in all specified test cases. These simulations combined with reasoning based on admissible degrees in Figure 1, led us to state conjectures concerning the solution to (71). For reproducibility, all simulations (including starting values) are provided in the publicly available GitHub repository: https://github.com/GustafLorentzon/polynomial-set-paper.
@@ 5.1 Four multiplications
When we study $m=4$, we identify from Figure 1 that the highest degree of the admissible polynomials is $12$, since without reductions we only have $16$ degrees of freedom, which cannot parameterize $\Pi_{16}$. The solution to (71) is indeed 12. This can already be concluded from the method in [24, p. 237] which is a method to evaluate polynomials of degree 12, given in their monomial basis, using only $m=4$ multiplications. In our terminology this corresponds to the reduction $a_{2,3}=0$ and, additionally, $a_{4,2}=0$. The method in [24, p. 237] involves square roots of expression containing the monomial coefficients, roots of a polynomial of degree four, as well as divisions of certain quantities leading to exceptions corresponding to some limit cases. Therefore, from [24] we conclude that
@@
We now present a slightly more general method for evaluating any polynomial in $\Pi_{12}$ with four multiplications. The new method does not involve square roots or divisions except for the leading monomial coefficient. Consider the evaluation schemes with the following structure
@@
[…]
@@
With the conditions $c_{2}=\alpha_{1}$, $c_{1}=\alpha_{0}$, $b_{43}=\beta_{43}-a_{43}$ and $b_{42}=\beta_{42}-a_{42}$, we have explicitly computed all variables in (73).
Recall that $\alpha_{12}\neq 0$ for $p\in\Pi_{12}$; therefore, we have made no assumptions other than the degree of the polynomial. Moreover, the formulas preserve the algebraic structure of the variables, e.g., if $\alpha_{0},\ldots,\alpha_{12}\in\mathbb{R}$, then $(A,B,c)$ is a real triplet, so the evaluation coefficients are real. We conclude that
@@
holds in both a real and a complex sense.
@@ 5.2 Five multiplications
For $m=5$ multiplications we see in Figure 1 that the highest degree of the admissible polynomials is $20$. To our knowledge, the state of the art is $d=18$ as given in [25, Equation (17)-(19)] with $s=2$ which is based on [24] combined with the Paterson–Stockmeyer evaluation. In our terminology, that approach corresponds to the reduction $a_{23}=a_{56}=0$ and additionally imposing $a_{42}=0$. The reduction in Figure 1 leading to a polynomial of degree $d=20$ corresponds to $a_{45}=a_{56}=0$.
We were not able to explicitly derive a solution to the multivariate polynomial system for the structure with $d=20$ analytically; instead, we needed to resort to computational tools. The Julia package HomotopyContinuation.jl [4] includes methods to solve polynomial systems of equations based on numerical continuation and with advanced initialization of starting points for the homotopy method. We implemented the system equations derived from considering each monomial coefficient for the data structures of this package. Since we have more variables than equations for this structure, we empirically fixed some variables, choosing which variable to fix by trying to reduce the total degree of the system as much as possible. We additionally solved those equations that could be solved explicitly, e.g., the first and last equations.
With this setup, we were able to compute $(A,B,c)$ for a large number of given polynomials $p\in\Pi_{20}$, including the truncated Taylor expansion of the exponential as well as the function 1/(1x)1/(1-x). The simulations were done in Julia, and the code is given in the GitHub repository. Moreover, code to actually evaluate polynomials is provided in both Julia and Matlab. For conciseness, we report only the numbers for the matrix exponential in the following.
In an attempt to prevent large condition numbers, we sought solutions that did not have excessively large or small values. In this case, we mitigated large numbers by scaling the input. For the exponential, we approximate $e^{\alpha x}$ with $\alpha=8$ fixed, since this makes $c_{7}=\alpha^{20}/20!\approx 0.47$ in the order of magnitude one. Although the input scaling can be reversed by transforming entries in the table, it was not deemed numerically useful and therefore it was not included in the following presentation of results.
HomotopyContinuation.jl found several solutions and the solution with the smallest values was the following
@@
The upcoming work [26] suggests that there is indeed a constructive way to form such evaluation schemes for five multiplications.